During a process a system releases 12.4 J of heat and does 4.2 J of work on the surroundings. The change in internal energy (ΔU) of the system is _____ J.

and the answer is not -8.2
Answer #1

First Law of Thermodynamics

Q-W = U

U = 12.4 -(-4.2) = 16.6 kJ